5.8 + [-log(10^-6)-log(1)-2](2.9)= x
Answer by nerdybill(7384)

5.8 + [-log(10^-6)-log(1)-2](2.9)= x
5.8 + [-(-6)-0-2](2.9)= x
5.8 + [6-0-2](2.9)= x
5.8 + [4](2.9)= x
5.8 + 11.6= x
5.8 + 11.6= x
17.4=x